# Kleene Algebrae, Kleene Modules, and Morita Equivalence

### v1

@@
Kleene Algebrae, Kleene Modules, and Morita Equivalence
Luke Serafin (lss255)
(October 28, 2025)
@@ 1 Introduction
Modules and the notion of Morita equivalence are foundational to the classical study of rings. These concepts extend naturally to semirings and then specialize to Kleene algebrae, and my goal is to investigate Kleene modules and Morita equivalence of Kleene algebrae in the hope that some of the power seen in the context of rings may be found in this new context as well. I undertook the definition and study of Kleene modules on my own, but was aware that others had done so and came across the notion in a form—which is exactly what I settled on—in [1]. I was not until recently aware of the fact that Morita equivalence has been studied in the context of semirings, but it has. See for example [2].
2 Kleene modules
@@
Let $K$ be a Kleene algebra. A left Kleene module over $K$ is a left idempotent semiring module $M$ over the semiring reduct of $K$ which additionally satisfies that $am\leq m$ implies $a^{*}m\leq m$ for $m\in M$ and $a\in K$. A right Kleene module is similarly a right module satisfying that $ma\leq m$ implies $ma^{*}\leq m$. A Kleene $(K,L)$-bimodule is a structure which is simultaneously a left Kleene $K$-module and a right Kleene $L$-module, such that if $m\in M$, $a\in K$, and $b\in L$, then $(am)b=a(mb)$.
Kleene modules are a quasivariety, and so by universal algebra there are free Kleene $K$-modules over arbitrary sets of generators for $K$ a fixed Kleene algebra.
Examples
Algebrae
Any Kleene algebra $K$ is a $(K,K)$-bimodule over itself, and a $(2,2)$-bimodule over the Kleene algebra $2$ with two elements. Moreover it is an $(A,A)$-bimodule over any subalgebra $A$ of $K$. In particular, if $K$ is a Kleene algebra with tests (a KAT), and $B$ is its distinguished Boolean subalgebra, then $B$ is a Kleene subalgebra of $K$ and $K$ is a $(B,B)$-bimodule.
Ideals
Let $I$ be a left ideal of a Kleene algebra $K$. Then $I$ is a left Kleene module with action given by multiplication. To see this, note that $I$ is clearly a left idempotent semiring module under the action of multiplication, and that for $m\in I$, if $am\leq m$ then by the *-axioms $a^{*}m\leq m$. Similary, if $J$ is a right ideal of $K$ then $J$ is a right Kleene module, and if $J$ is a two-sided ideal then it is a $(K,K)$-Kleene bimodule.
Quotients
Let $\equiv$ be a congruence of a Kleene algebra $K$, and set $Q=K/{\equiv}$. Then $Q$ is a Kleene $K$-bimodule with action given by multiplication. This follows from the fact that $\equiv$ is a congruence, and provides an example of a Kleene algebra with a natural Kleene bimodule structure.
Function modules
Let $K$ be a Kleene algebra, $B$ be a set, and consider the set $K^{B}$ of functions from $B$ to $K$ endowed with pointwise addition and zero as the constant-zero function. Define left and right actions of $K$ on $K^{B}$ by $(am)(b)=am(b)$ and $(ma)(b)=m(b)a$ for $a\in K$, $m\in K^{B}$. It is immediate that these operations give a $(K,K)$-bimodule structure on $K^{B}$. Moreover, if $B$ is finite, then $K^{B}$ has the usual finite basis given by characteristic functions for singletons in $B$, and is the free $(K,K)$-bimodule with generators $B$. Forgetting either the right or the left action, it is also the free left and the free right $K$-module with generators $B$. One can also restrict from $K^{B}$ to functions supported on a set in some lattice-ideal for $\operatorname{\mathcal{P}}(B)$. It is easy to see that the finitely-supported elements of $K^{B}$ form the free $K$-module with generators $B$.
Homomorphism modules
Fix a Kleene algebra $K$ and let $M$, NN be left $K$-modules. The set $\operatorname{Hom}(M,N)$ becomes a left $K$-module when endowed with pointwise operations. Similarly for $M$, NN right $K$-modules, or $(K,S)$-bimodules where $S$ is another Kleene algebra.
Endomorphism modules
Let $x\in K$, $\alpha\in\mathop{\text{End}}(M)$. If $x\alpha\leq\alpha$, then for every $k\in K$, $x\alpha(k)\leq\alpha(k)$, so $x^{*}\alpha(k)\leq\alpha(k)$ for every $k\in K$ and hence $x^{*}\alpha\leq\alpha$. Consequently $\mathop{\text{End}}(K)$ is a left Kleene module under multiplication by elements of $K$. Similarly it is a right Kleene module under multiplication by elements of $K$, and it is easy to check that it is a $(K,K)$-bimodule. Matrices are the special case of this where $M=K^{n}$.
Propositional Dynamic Logic
Fix sets $P$ of action atoms and $T$ of test atoms, respectively. Let $K$ be the free Kleene algebra generated by $P$, and $B$ the free Boolean algebra generated by $T$. The Lindenbaum-Tarski algebra $A$ of propositional dynamic logic with action atoms $P$ and test atoms $T$ can be viewed as a $(B,B)$-bimodule via the subalgebra bimodule structure, but also has the structure of a left $K$-module via the action $pa=[\langle p\rangle\phi]$ where $p\in K$, $a\in A$, $a=[\phi]$, and brackets denote the equivalence class of a formula in the Lindenbaum-Tarski algebra. This follows immediately from the axioms for propositional dynamic logic.
The dual of a Kleene module
@@
Fix a Kleene algebra $K$, and let $M$ be a left (respectively, right) Kleene module over $K$. The dual $M^{\circ}$ of $M$ is the set of Kleene $K$-module homomorphisms from $M$ to $K$, with pointwise addition and 0 the constant 0-function.
The proof that the dual of a Kleene module is a Kleene module is immediate.
@@
For $M$ a left (respectively, right) Kleene module over $K$, its dual $M^{\circ}$ is a right (respectively, left) Kleene module over $K$. Moreover the dual of a Kleene $(K,S)$-bimodule is a Kleene $(S,K)$-bimodule.
@@
The dual of $M$ is isomorphic to the homomorphism module $\operatorname{Hom}(M,K)$. ∎
Tensor products of Kleene modules

### v2

# Morita Rigidity for Kleene Algebras

@@
Abstract
We introduce Morita equivalence to the study of Kleene algebras and modules. Classical characterizations of Morita-equivalent semirings such as having equivalent categories of modules and one semiring being a full matrix algebra over the other carry over. We also observe that Morita equivalence can be applied to extending and restricting scalars in Lindenbaum–Tarski algebras of propositional dynamic logics. But the signature result which we obtain is a form of rigidity for Kleene algebras, which states that if the semiring reducts of two Kleene algebras are Morita-equivalent, then the Morita equivalence is in fact witnessed by Kleene bimodules.
@@ 1 Introduction
Modules and the notion of Morita equivalence are foundational to the classical study of rings [15], and we assume the reader has some familiarity with the classical theory of modules though not necessarily of Morita equivalence.
The theory of modules has previously been extended to the context of semirings (for example see [5]), though we spend some time developing the theory of tensor products since these do not appear to have been previously worked out for Kleene modules, and the standard definition in [5] is inadequate for us because it makes all tensor products of modules over additively idempotent semirings (such as Kleene algebras) trivial. The alternative tensor product defined in [8] does have the properties we need, but the author had worked out what turned out to be the same definition as Katsov’s before discovering Katsov’s paper and so has included a full development of tensor products for Kleene algebras rather than simply extending Katsov’s definition for modules over semirings to the case of Kleene modules.
Additionally, a variety of instances of Kleene modules are presented, the most important being endomorphism modules of Kleene modules (including $n\times n$ matrices) and the Lindenbaum–Tarski algebra of propositional dynamic logic. Morita equivalence for rings is obtained by replacing the morphisms of the category of rings by bimodules, and considering the induced notion of isomorphism. More specifically, the morphisms from a ring $R$ to a ring $S$ are the $(R,S)$-bimodules, and composition of bimodules is given by tensor product. The author has undertaken to study Morita equivalence in the context of Kleene algebras, leading to the discovery of a number of results and directions for further study. Several standard results carry over, such as characterizations of Morita equivalence in terms of categorical equivalence of module categories, and in terms of matrix algebras defined in terms of full (multiplicative) idempotents. One surprising fact is that if the semiring reducts of two Kleene algebras $K$ and $S$ are Morita-equivalent, then in fact $K$ and $S$ are Morita-equivalent as Kleene algebras (using Kleene bimodules to witness the equivalence). This is a rigidity theorem, saying that a weaker equivalence (Morita equivalence in the category of semirings) implies a stronger equivalence (Morita equivalence in the category of Kleene algebras) within a particular domain of structures (here, Kleene algebras).
2 Semiring Preliminaries
We review here the basics of semirings and the corresponding generalization of module (sometimes called a semimodule in the literature). For more concerning semirings see for example [5] and [6]. In [5] the author develops a theory of tensor products of semirings, but these tensor products are always cancellative, which means trivial in the context of additively idempotent semirings with which we are concerned. Therefore we develop an alternative, and in the author’s opinion more natural, tensor product which is useful even in the case of additively idempotent semirings. Though this tensor product appears to be new, a form of tensor product for Kleene algebras is developed in [7], and the book [4] considers tensor products for a variety of idempotent structures related to Kleene algebras.
@@ 2 Semiring Preliminaries
A (unital) semiring is an algebraic structure $\langle S,+,\cdot,0,1\rangle$ such that
  1. 1.
    $\langle S,+,0\rangle$ is a commutative monoid,
  2. 2.
    $\langle S,\cdot,1\rangle$ is a monoid (not necessarily commutative),
  3. 3.
    $s\cdot 0=0\cdot s=0$ for every $s\in S$, and
  4. 4.
    multiplication distributes over addition, meaning that for any $x,y,z\in S$, $x\cdot(y+z)=(x\cdot y)+(x\cdot z)$ and $(x+y)\cdot z=(x\cdot z)+(y\cdot z)$.
We often abuse notation and denote the semiring $\langle S,+,\cdot,0,1\rangle$ by $S$, the operations being understood. Also, we generally denote multiplication $\cdot$ by juxtaposition, $xy=x\cdot y$, and regard multiplication as taking precedence over addition in expressions, so $xy+z$ is parsed as $(x\cdot y)+z$. When multiple semirings are under consideration we may add subscripts to operations, e.g. $\alpha(r+_{S}s)\cdot_{T}t$.
As is common in ring theory, we sometimes need multiplication in a semiring to be computed in the opposite order, and so define the operation on semirings which reverses the order of multiplication.
@@
For $S$ a semiring, the opposite semiring ${S}^{\mathrm{op}}$ has the same underlying set, addition, and constants as $S$, but the multiplication $\cdot_{{S}^{\mathrm{op}}}$ is defined by $s\cdot_{{S}^{\mathrm{op}}}t=t\cdot_{S}s$ for $s,t\in S$.
Note that $({S}^{\mathrm{op}})^{\mathrm{op}}=S$ for any semiring $S$. We next give the definition of a module over a semiring (also called a semimodule in the literature; e.g. [5]).
Definition 3.
Let $S$ be a semiring. A left semiring module over $S$ (or simply a left $S$-module, where it is understood that $S$ is a semiring) is a commutative monoid $\langle M,+,0\rangle$ together with a map $S\times M\rightarrow M$, denoted by juxtaposition—so $(s,m)\mapsto sm$ for $s\in S$ and $m\in M$—satisfying the following for any $s,t\in S$ and $m,n\in M$:
  1. 1.
    $s(m+n)=sm+sn$
  2. 2.
    $(s+t)m=sm+tm$
  3. 3.
    $s(tm)=(st)m$
  4. 4.
    $1m=m$
  5. 5.
    $0m=0$.
Right $S$-modules are defined symmetrically, with the operation now a map $M\times S\rightarrow M$, written on the right: $(m,s)\mapsto ms$.
For $S$ and $T$ semirings, an $(S,T)$-bimodule is a commutative monoid $\langle M,+,0\rangle$ equipped with both a left $S$-module map and a right $T$-module map, such that these two structures are compatible in the sense that for any $s\in S$, $t\in T$, and $m\in M$, $s(mt)=(sm)t$. As in the case of semirings, when we need to disambiguate in which module an operation is happening, we use subscripts, as in $s\cdot_{N}\alpha(m+_{M}n\cdot_{M}t)$ where $M$ is a right semiring module and $N$ a left semiring module over a semiring $S$. As in the literature on modules over rings, we sometimes use subscripts to the left or right of a symbol denoting a module over a semiring to indicate both which semiring acts on the module and from which side. Thus if $S$, $T$ are semirings, ${{}_{S}}M$ denotes a left module over $S$, $N_{T}$ denotes a right module over $T$, and ${{}_{S}}B_{T}$ indicates an $(S,T)$-bimodule. This notation is particularly helpful when working with direct sums and tensor products, as we shall later.
To define tensor products we shall need the notion of a congruence relation. This is a general notion from universal algebra (see for example [3]). We view modules as algebraic structures with a potentially infinite number of unary operations of scalar multiplication, one for each scalar.
Definition 4.
A congruence on an algebraic structure $\mathfrak{A}$ is an equivalence relation on the underlying set of $\mathfrak{A}$ which preserves all operations. For example, if $\mathfrak{A}$ is a semiring and $\equiv$ is a congruence on $\mathfrak{A}$, then for $a,b,a^{\prime},b^{\prime}$ elements of the underlying set of $\mathfrak{A}$, if $a\equiv a^{\prime}$ and $b\equiv b^{\prime}$, then $a+b\equiv a^{\prime}+b^{\prime}$ and $ab\equiv a^{\prime}b^{\prime}$.
It is a well-known fact from universal algebra (see [3] for a proof) that if $\equiv$ is a congruence on an algebraic structure $\mathfrak{A}$, then the set of equivalence classes of $\equiv$ carry a structure of the same type as $\mathfrak{A}$, with operations computed by taking the equivalence class of the corresponding operation in $\mathfrak{A}$ performed on representatives. This is the quotient of the algebra $\mathfrak{A}$ by the congruence $\equiv$. Thus for example if $\equiv$ is a congruence on a semiring $S$ and we use brackets to denote $\equiv$-equivalence classes, then in the quotient of $S$ by $\equiv$ we have for $s,t\in S$ that $[s]+[t]=[s+t]$ and $[s][t]=[st]$. It is easy to use the stipulation that $\equiv$ is a congruence to prove that these operations are well-defined.
As in the case of semirings, we can define the opposite of a semiring module.
Definition 5.
For $S$ a semiring and $M$ a left semiring module over $S$, the opposite module ${M}^{\mathrm{op}}$ to $M$ is a right ${S}^{\mathrm{op}}$-module with the action of ${S}^{\mathrm{op}}$ given by $m\cdot_{{M}^{\mathrm{op}}}s=s\cdot_{M}m$ for $s\in S$ and $m\in M$. The opposite of a right $S$-module and the opposite of an $(S,T)$-bimodule for $T$ also a semiring are defined similarly, with the opposite of a right $S$-module being a left ${S}^{\mathrm{op}}$-module and the opposite of an $(S,T)$-bimodule being a $({T}^{\mathrm{op}},{S}^{\mathrm{op}})$-bimodule.
@@
The opposite of a module over a semiring $S$ is a module of the appropriate type.
@@
Let $S$ and $T$ be semirings; we treat just the case of an $(S,T)$-bimodule, since the other cases are straightforward adaptations of the argument. Fix an $(S,T)$-bimodule $M$ and note that the actions of ${T}^{\mathrm{op}}$ and of ${S}^{\mathrm{op}}$ on ${M}^{\mathrm{op}}$ are given respectively by
$t\cdot_{{M}^{\mathrm{op}}}m=m\cdot_{M}t\quad\text{and}\quad m\cdot_{{M}^{\mathrm{op}}}s=s\cdot_{M}m$
for $s\in S$, $t\in T$, and $m\in M$. It is immediate that these actions are additive, preserve zero, and have the property that $1_{{S}^{\mathrm{op}}}$ and $1_{{T}^{\mathrm{op}}}$ act as identities. It remains to verify that for $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, and $m\in M$, we have the following:
$t\cdot_{{M}^{\mathrm{op}}}(t^{\prime}\cdot_{{M}^{\mathrm{op}}}m)=(t\cdot_{{M}^{\mathrm{op}}}t^{\prime})\cdot_{{M}^{\mathrm{op}}}m,$ (1)
$(m\cdot_{{M}^{\mathrm{op}}}s)\cdot_{{M}^{\mathrm{op}}}s^{\prime})=m\cdot_{{M}^{\mathrm{op}}}(s\cdot_{{M}^{\mathrm{op}}}s^{\prime}),\quad\text{and}$ (2)
$(t\cdot_{{M}^{\mathrm{op}}}m)\cdot_{{M}^{\mathrm{op}}}s=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{{M}^{\mathrm{op}}}s).$ (3)
For ((1)) we compute that
$t\cdot_{{M}^{\mathrm{op}}}(t^{\prime}\cdot_{{M}^{\mathrm{op}}}m)$ $=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{M}t^{\prime})=(m\cdot_{M}t^{\prime})\cdot_{M}t=m\cdot_{M}(t^{\prime}\cdot_{T}t)=m\cdot_{M}(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})=$
$=(t\cdot_{{T}^{\mathrm{op}}}t^{\prime})\cdot_{{M}^{\mathrm{op}}}m.$
The case of ((2)) is symmetric. For ((3) note that
$(t\cdot_{{M}^{\mathrm{op}}}m)\cdot_{{M}^{\mathrm{op}}}s=(mt)\cdot_{{M}^{\mathrm{op}}}s=s(mt)=(sm)t=t\cdot_{{M}^{\mathrm{op}}}(sm)=t\cdot_{{M}^{\mathrm{op}}}(m\cdot_{{M}^{\mathrm{op}}}s).$
Proposition 7.
For $S$ a semiring, ${S}^{\mathrm{op}}$ has the structure of an $(S,S)$-bimodule (and hence also of either a left or a right semiring module over $S$).
Proof.
To avoid confusion, we shall denote multiplication in $S$ by $\cdot_{S}$, multiplication in ${S}^{\mathrm{op}}$ by $\cdot_{{S}^{\mathrm{op}}}$, the left action of $S$ on ${S}^{\mathrm{op}}$ by $\cdot_{L}$, and the right action of $S$ on ${S}^{\mathrm{op}}$ by $\cdot_{R}$. Given $a,b\in S$, define the left and right actions of $S$ by
$a\cdot_{L}b=a\cdot_{R}b=b\cdot_{{S}^{\mathrm{op}}}a.$
Because the two actions are given by the same operation, the fact that they commute follows immediately from associativity of multiplication in ${S}^{\mathrm{op}}$. We check that multiplication in $S$ distributes over the right action of $S$; the case of the left action is symmetric. For $a,b,c\in S$, compute
$a\cdot_{R}(b\cdot_{S}c)=(b\cdot_{S}c)\cdot_{{S}^{\mathrm{op}}}a=(c\cdot_{{S}^{\mathrm{op}}}b)\cdot_{{S}^{\mathrm{op}}}a=c\cdot_{{S}^{\mathrm{op}}}(b\cdot_{{S}^{\mathrm{op}}}a)=c\cdot_{R}(b\cdot_{R}a).$
3 Kleene Algebras and Kleene Modules
A semiring $S$ is additively idempotent if and only if for every $\in S, x+x$=xx+x=x. Additively idempotent semirings (and in fact idempotent commutative monoids more generally) [^1] Commutative idempotent monoids are easily seen to be definitionally equivalent to join-semilattices come equipped with a partial order defined by $x\leq y$ if and only if $x+y=y$. It is easy to check that $x+y$ is then the least upper bound of $x$ and $y$ in the order.
@@ 3 Kleene Algebras and Kleene Modules
A Kleene algebra $\langle K,+,\cdot,0,1,*\rangle$ is an additively idempotent semiring with an additional unary operation * that satisfies the following axioms for any $x,y,z\in K$:
  • $1+xx^{*}=1+x^{*}x=x^{*}$,
  • if $x+yz\leq z$ then $y^{*}x\leq z$,
  • if $x+zy\leq z$ then $xy^{*}\leq z$.
Note that, since semirings are defined to be unital, if a semiring $S$ is additively idempotent then any left or right module over $S$ is also additively idempotent. The argument for a left module $M$ over $S$ will be sufficient, because the case of a right module is symmetric. For $m\in M$,
$m+m=1m+1m=(1+1)m=1m=m.$
Also note that by the symmetry of the conditions on the asterate, the opposite semiring of a Kleene algebra becomes a Kleene algebra when endowed with the same asterate.
Definition 9.
Fix a semiring $S$, and let $M$ and $N$ be left $S$-modules. A map $\phi:M\rightarrow N$ is a homomorphism of left $S$-modules if and only if for every $s\in S$ and $m,n\in M$,
$\phi(m+n)=\phi(m)+\phi(n)\quad\text{and}\quad\phi(sm)=s\phi(m);$
the definitions of homomorphisms of right $S$-modules and of $(S,T)$-bimodules for $T$ a semiring are analogous. The set of all $S$-module homomorphisms from $M$ to $N$ is denoted $\operatorname{Hom}(M,N)$. Note that this set naturally comes equipped with pointwise addition; later we shall see how it can be endowed with module structure when $M$ and $N$ have additional structure. When $M=N$ we write $\mathop{\text{End}}(M)$ for $\operatorname{Hom}(M,M)$ and speak of endomorphisms of the semiring module $M$. The category of left $S$-modules and homomorphisms [^2] With functional composition as the composition operation. is denoted $\text{Mod}_{S}$, the category of right $S$-modules and homomorphisms is denoted $_{S}\text{Mod}$}, and the category of $(S,T)$-bimodules and homomorphisms is denoted $_{S}\text{Mod}_{T}$}. In all these categories isomorphism has its usual, categorical meaning (for which see e.g. [16]).
Definition 10.
Let $K$ be a Kleene algebra. A left Kleene module over $K$ is a semiring module $M$ over the semiring reduct of $K$ satisfying the following implication for $m\in M$ and $a\in K$:
$am\leq m\Rightarrow a^{*}m\leq m.$
A right Kleene module over $K$ is similarly a right semiring module $M$ satisfying $ma\leq m\Rightarrow ma^{*}\leq m$ for $m\in M$ and $a\in K$. A Kleene $(K,L)$-bimodule is a semiring $(K,L)$-bimodule which is simultaneously a left Kleene module over $K$ and a right Kleene module over $L$.
Kleene modules are a quasivariety (see e.g. [10], which gives a universal Horn axiomatization), and so by universal algebra there are free Kleene $K$-modules over arbitrary sets of generators for $K$ a fixed Kleene algebra (see for example [1]).
3.1 Examples
Algebras. Any Kleene algebra $K$ is a $(K,K)$-bimodule over itself, and a $(2,2)$-bimodule over the Kleene algebra $2$ with two elements. Moreover it is an $(A,A)$-bimodule over any subalgebra $A$ of $K$. In particular, if $K$ is a Kleene algebra with tests (as defined in [14]; a KAT), and $B$ is its distinguished Boolean subalgebra, then $B$ is a Kleene subalgebra of $K$ and $K$ is a $(B,B)$-bimodule.
Ideals. Let $I$ be a left ideal of a Kleene algebra $K$. Then $I$ is a left Kleene module with action given by multiplication. To see this, note that $I$ is clearly a left module over an additively idempotent semiring under the action of multiplication, and that for $m\in I$, if $am\leq m$ then by the *-axioms for Kleene algebras, $a^{*}m\leq m$. Similary, if $J$ is a right ideal of $K$ then $J$ is a right Kleene module, and if $J$ is a two-sided ideal then it is a Kleene $(K,K)$-bimodule.
Quotients. Let $\equiv$ be a congruence of a Kleene algebra $K$, and set $Q=K/{\equiv}$. If $Q$ is in fact a Kleene algebra, then $Q$ is a Kleene $K$-bimodule with action given by multiplication. This follows from the fact that $\equiv$ is a congruence, and provides an example of a Kleene algebra construction that comes with a natural Kleene bimodule structure.
Function modules. Let $K$ be a Kleene algebra, $B$ be a set, and consider the set $K^{B}$ of functions from $B$ to $K$ endowed with pointwise addition and zero as the constant-zero function. Define left and right actions of $K$ on $K^{B}$ by $(am)(b)=am(b)$ and $(ma)(b)=m(b)a$ for $a\in K$, $m\in K^{B}$. It is immediate that these operations give a $(K,K)$-bimodule structure on $K^{B}$. Moreover, if $B$ is finite, then $K^{B}$ has the usual finite basis given by characteristic functions for singletons in $B$, and hence is the free $(K,K)$-bimodule with generators $B$. Forgetting either the right or the left action, it is also the free left and the free right $K$-module with generators $B$. One can also restrict from $K^{B}$ to functions supported on a set in some lattice-ideal for $\operatorname{\mathcal{P}}(B)$. It is easy to see that the finitely-supported [^3] So, supported in the ideal of finite sets. elements of $K^{B}$ form the free $K$-module with generators $B$.
Homomorphism and endomorphsim modules. Fix Kleene algebras $K$ and $S$ and let $M$ and $N$ be $(K,S)$-bimodules. Denote by $H$ the set $\operatorname{Hom}(M,N)$, which we shall soon endow with additional structure. For $\alpha,\beta\in\operatorname{Hom}(M,N)$, define $\alpha+\beta$ to be the pointwise sum of $\alpha$ and $\beta$ and $0_{H}$ to be the constant function with value zero. This trivially gives $\operatorname{Hom}(M,N)$ the structure of a commutative monoid. Now define a left action of $S$ by $(s\cdot_{H}\alpha)(m)=\alpha(m\cdot_{M}s)$ for $s\in S$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$; it is straightforward to check that this gives $\operatorname{Hom}(M,N)$ the structure of a left module over $S$. It remains to check for $s\in S$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$ that if $s\alpha\leq\alpha$ then $s^{*}\alpha\leq\alpha$. Assume that $s\alpha\leq\alpha$, so for each $m\in M$, $s\alpha(m)\leq\alpha(m)$. Applying the action gives that $\alpha(m\cdot_{M}s)=\alpha(m)\cdot_{N}s\leq\alpha(m)$, so because $M$ is a right semiring module over $S$, $\alpha(m)\cdot_{N}s^{*}=\alpha(m\cdot_{M}s^{*})\leq\alpha(m)$, which means $s^{*}\cdot_{H}\alpha(m)\leq\alpha(m)$, that is, $s^{*}\cdot_{H}\alpha\leq\alpha$.
Reversing order when appropriate handles the case of homomorphisms of into right modules Specifically, if $M$ and $N$ are $(S,K)$-bimodules, then $\operatorname{Hom}(M,N)$ becomes a right $K$-module when endowed with pointwise addition, the constant function zero, and the right action of $K$ defined by $(\alpha\cdot_{H}k)(m)=\alpha(k\cdot_{M}m)$ for $k\in K$, $m\in M$, and $\alpha\in\operatorname{Hom}(M,N)$.
Let $K$ be a Kleene algebra, $M$ a $(K,K)$-bimodule. Define
$\mathop{\text{End}}(M)=\operatorname{Hom}(M,M),$
the endomorphisms of $M$. By the preceding example $\mathop{\text{End}}(M)$ is a $(K,K)$-bimodule over the semiring reduct of $K$. Now we verify that it in fact carries the structure of a Kleene $(K,K)$-bimodule. For $k\in K$ and $\alpha\in\mathop{\text{End}}(M)$, if $\alpha k\leq\alpha$, then for every $m\in M$, $\alpha(m)k\leq\alpha(m)$, so $\alpha(m)k^{*}\leq\alpha(m)$ for every $m\in M$ and hence $\alpha k^{*}\leq\alpha$. Consequently $\mathop{\text{End}}(K)$ is a right Kleene module over $K$. A similar argument gives that $\mathop{\text{End}}(K)$ is a left Kleene module over $K$, so it remains to verify that the left and right actions of $K$ commute. This is given by the following computation, where $s,k\in K$, $m\in M$, and $\alpha\in\mathop{\text{End}}(K)$.
$(s\alpha)k](m)=(s\alpha)(km)=\alpha(kms)=(\alpha k)(ms)=[s(\alpha k)](m).$
In the special case of $K^{n}$ viewed as a $({K}^{\mathrm{op}},{K}^{\mathrm{op}})$-bimodule with pointwise scalar multiplication, ${\mathop{\text{End}}(M)}^{\mathrm{op}}$ is the $(K,K)$-bimodule of $n\times n$ matrices over $K$. The reason for taking the opposite is that the left and right actions of $K$ on $M$ are reversed in the definition of $\operatorname{Hom}(M,M)=\mathop{\text{End}}(K)$.
Also, as in the classical theory of modules over rings, the dual $M^{\circ}$ of a Kleene $(K,K)$-bimodule $M$ is the Kleene $(K,K)$-bimodule $\operatorname{Hom}(M,{{}_{K}}K_{K})$. Since the concepts of homomorphism module, endomorphism module, and the dual of a module will be used below, we recapitulate the definitions here.
Definition 11.
Let $K$, $S$ be Kleene algebras and $M$, NN Klene $(K,S)$-bimodules. The homomorphsim module is the set of homomorphisms from $M$ to $N$, endowed with the operations defined at the beginning of this example. If $K=S$, the endomorphism module of $M$ is $\mathop{\text{End}}(M)=\operatorname{Hom}(M,M)$, and the dual of $M$ is $M^{\circ}=\operatorname{Hom}(M,{{}_{K}}K_{K})$.
Propositional Dynamic Logic. Fix sets $P$ of action atoms and $T$ of test atoms, respectively. [^4] For details concerning propositional dynamic logic see for instance [13]. Let $K$ be the free Kleene algebra generated by $P$, and $B$ the free Boolean algebra generated by $T$. The Lindenbaum-Tarski algebra $A$ of propositional dynamic logic with action atoms $P$ and test atoms $T$ can be viewed as a $(B,B)$-bimodule via the subalgebra bimodule structure, but also has the structure of a left $K$-module via the action $pa=[\langle p\rangle\phi]$ where $p\in K$, $a\in A$, $a=[\phi]$, and brackets denote the equivalence class of a formula in the Lindenbaum-Tarski algebra. This follows immediately from the axioms for propositional dynamic logic.
3.2 Tensor products of Kleene modules
Tensor products of modules are defined so that the tensor product functor $\otimes N$ is left-adjoint to the hom-functor $\operatorname{Hom}(N,{-})$ for each module $N$ of the correct type. This is constructed by taking the quotient of formal sums of pairs from ${{}_{A}}M_{B}\times{{}_{B}}N_{C}$ by the least congruence imposing $A$-linearity in the first coordinate, $C$-linearity in the second coordinate, and balance, which means that for any $m\in M$, $n\in N$, and $b\in B$, $(mb,n)$ is identified with $(m,bn)$. But in the context of Morita equivalence only the isomorphism class of tensor products is relevant, so we define what it means to be a tensor product via a categorical adjunction requirement and then check that the definition just outlined satisfies this requirement.
Definition 12.